how do you factor x^3-4x^2-5x+20
Answer by stanbon(75887)

factor x^3-4x^2-5x+20
---------
I graphed it and found a zero at x = 4.
---
Use synthetic division to find the other roots.
---
4)....1....-4....-5....20
.......1.....0.....-5...|..0
Quotient x^2-5
Remainder : zero
-----------------------
Solve x^2-5 = 0
x = sqrt(5) or x = -sqrt(5)
------------------------------------
The three roots are 4, sqrt(5), -sqrt(5)
=========================================
Cheers,
Stan H.